# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 8. Conclusion and Future Work
We have studied the complexity of resilience for RPQs, towards the goal of characterizing the languages for which the problem is tractable. We study bag and set semantics, but our results show no difference between the two settings, unlike for CQs (makhija2024unified). We leave open the question of whether there are languages $L$, regular or not, for which $\mathtt{RES}_{\mathrm{set}}(L)$ is PTIME but $\mathtt{RES}_{\mathrm{bag}}(L)$ is NP-hard. We also leave open the question of classifying the complexity of resilience in intermediate settings, e.g., bag semantics where only weights $1$, and $\infty$ are allowed; or the setting where some relations are indicated as exogenous in the signature, amounting to giving a weight of $\infty$ to all their facts.
Our work does not fully characterize the tractability boundary and leaves some cases open. At a high level, we still do not understand the non-local languages that are not four-legged, i.e., non-letter-Cartesian languages featuring counterexamples where some of the “legs” are empty. We also do not know if the submodularity approach is needed for queries like $abc|be$, or whether tractability can be obtained in easier ways. [^6] After we had submitted this paper for review, an argument was pointed out by Martín Muñoz (cstheory_martin) implying that Section 7.4 can be shown using a flow encoding instead of submodular function minimization. Further, his argument seems to imply that the tractability of resilience extends to more queries, such as $abcd|be$. We leave the exploration of these consequences to a future version of the paper.
@@
Another open question concerns non-Boolean RPQs, i.e., where endpoints are fixed (or given as input). For instance, the RPQ $aa$ then seems tractable as possible contingency sets have a simple structure; but longer RPQs may stay hard. This problem may relate to length-bounded cuts (itai1982complexity; 10.1145/1868237.1868241), or multicuts (garg1994multiway) and skew multicuts (kratsch2015fixed). Note that this non-Boolean setting corresponds to the problem of deletion propagation with source side effects (buneman2002propagation), which was studied before resilience: this problem asks about finding a minimum contingency set to remove a specific query answer. The case of removing multiple tuples has also been investigated (kimelfeld2013multituple). We do not know whether a dichotomy in one of these settings implies a dichotomy for our problem, or vice-versa. We also do not understand how our results relate to VCSPs (bodirsky2024complexity), e.g., whether classes of VCSPs with better algorithms (in combined complexity, or in terms of the polynomial degree) can imply tractable algorithms for some RPQ classes; or whether VCSP dichotomies hold in set semantics (or in bag semantics with weights restricted to $1,\infty$).
@@
Last, resilience can be studied for recursive queries beyond RPQs, in particular two-way RPQs or Datalog, though this would require new techniques (unlike RPQs, these queries are not “directional”).
@@ 9. Acknowledgments
The authors are grateful to Charles Paperman for help about local languages, and to the reviewers for their valuable feedback. Part of this work was conducted while the authors were visiting the Simons Institute for the Theory of Computing. Amarilli was partially supported by the ANR project EQUUS ANR-19-CE48-0019, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 431183758, and by the ANR project ANR-18-CE23-0003-02 (“CQFD”). Gatterbauer and Makhija were partially supported by the National Science Foundation (NSF) under award number IIS-1762268.

### v3

@@ 8. Conclusion and Future Work
We have studied the complexity of resilience for RPQs, towards the goal of characterizing the languages for which the problem is tractable. We study bag and set semantics, but our results so far show no difference between the two settings, unlike for CQs (makhija2024unified, ). Refer back to Figure 1 on page 1 for a summary of our results and of some remaining open cases. We leave open the general question of whether there are languages $L$, regular or not, for which $\mathtt{RES}_{\mathrm{set}}(L)$ is PTIME but $\mathtt{RES}_{\mathrm{bag}}(L)$ is NP-hard. We also leave open the question of classifying the complexity of resilience in intermediate settings, e.g., bag semantics where only weights $1$ and $\infty$ are allowed; or the setting where some relations are indicated as exogenous in the signature, which amounts to giving a weight of $\infty$ to all their facts. Other settings left open by our work are those where we delete vertices (or just vertex labels) instead of edges, or where vertices and/or edges can be labeled by words or by non-empty words instead of single letters.
Our work does not fully characterize the tractability boundary and leaves some cases open. In particular, at a high level, we still do not fully understand the non-local languages that are not four-legged, i.e., non-letter-Cartesian languages featuring counterexamples where some of the “legs” are empty.
@@
Another open question concerns non-Boolean RPQs, i.e., where endpoints are fixed (or given as input). For instance, the RPQ $aa$ then seems tractable as possible contingency sets have a simple structure; but longer RPQs may stay hard. This problem may relate to length-bounded cuts (itai1982complexity, ; 10.1145/1868237.1868241, ), or multicuts (garg1994multiway, ) and skew multicuts (kratsch2015fixed, ). Note that this non-Boolean setting corresponds to the problem of deletion propagation with source side effects (buneman2002propagation, ), which was studied before resilience: this problem asks about finding a minimum contingency set to remove a specific query answer. The case of removing multiple tuples has also been investigated (kimelfeld2013multituple, ). We do not know whether a dichotomy in one of these settings implies a dichotomy for our problem, or vice-versa. We also do not understand how our results relate to VCSPs (bodirsky2024complexity, ), e.g., whether classes of VCSPs with better algorithms (in combined complexity, or in terms of the polynomial degree) can imply tractable algorithms for some RPQ classes; or whether VCSP dichotomies hold in set semantics (or in bag semantics with weights restricted to $1$ and $\infty$).
@@
Last, resilience can be studied for recursive queries beyond RPQs, in particular two-way RPQs or Datalog, though this would require new techniques (unlike RPQs, these queries are not “directional”). We note that for two-way RPQs, and more generally for queries expressible as monadic Datalog, the existence of finite duals implies a dichotomy on the resilience problem in bag semantics via the VCSP techniques of (bodirsky2024complexity, ).
@@ 9. Acknowledgments
The authors are grateful to Charles Paperman for help about local languages, to Rémi de Pretto for pointing out a minor problem, to Blaz Korecic and Martín Andrighetti for insightful discussions towards the proof of Proposition 7.9, and to the reviewers of the conference version for their valuable feedback. Part of this work was conducted while four of the authors were participating in the semester-long program Logic and Algorithms in Database Theory and AI at the Simons Institute for the Theory of Computing. Amarilli was partially supported by the ANR project EQUUS ANR-19-CE48-0019, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 431183758, and by the ANR project ANR-18-CE23-0003-02 (“CQFD”). Gatterbauer and Makhija were partially supported by the National Science Foundation (NSF) under award number IIS-1762268.